1.  Consider the following joint probability density function (pdf) of the random variables X and Y:   f(x,y)  =  0 ≤ x ≤ 6,    0 ≤ y ≤ 8     =  0  elsewhere                                                                           

Calculate the constant c if P(2 < X < c, 3 < Y < 5) = 0,06                             


1
Expert's answer
2021-07-19T12:06:23-0400
"f(x, y)=\begin{cases}
 1/48 & 0 ≤ x ≤ 6, 0 ≤ y ≤ 8 \\
 0 &elsewhere
\end{cases}"

Check


"\displaystyle\int_{-\infin}^{{\infin}}\displaystyle\int_{-\infin}^{\infin}f(x, y)dydx=\displaystyle\int_{0}^{{6}}\displaystyle\int_{0}^{8}\dfrac{1}{48}dydx"

"=\dfrac{1}{48}\displaystyle\int_{0}^{{6}}[y]\begin{matrix}
 8\\
 0
\end{matrix}dx=\dfrac{1}{6}[x]\begin{matrix}
 6\\
 0
\end{matrix}=1"



"P(2<X<c, 3<Y<5)=\displaystyle\int_{3}^{5}\displaystyle\int_{2}^{c}\dfrac{1}{48}dxdy"

"=\dfrac{1}{48}\displaystyle\int_{3}^{5}[x]\begin{matrix}
 c \\
 2
\end{matrix}dy=\dfrac{c-2}{48}[y]\begin{matrix}
 5 \\
 3
\end{matrix}=\dfrac{c-2}{24}=0.06"

"c=3.44"